OpenStudy (anonymous):

Integrate (6x+5)/(x^2+4x+4) from 1 to 9 is A/denominator + B/denominator

OpenStudy (anonymous):

So you need to do partial fractions for this. So, (6x+5)/(x+2)(x+2) = A/x+2 + B/(x+2)^2

OpenStudy (anonymous):

Solve for A and B, Ax^2+4Ax+4A + Bx+2B = 6x+5 4A+B=6 4A+2B=5 A= 7/4 B=-1

OpenStudy (anonymous):

Then find \[\int\limits_{}^{}(7/4)/(x+2) + \int\limits_{}^{}-1/(x+2)^2\]

OpenStudy (anonymous):

I'll just skip to the answer for now, You should wind up with 6ln(11/3) - 56/33 = 6.09873